A clinic employs nine physicians. Five of the physicians are female. Four patients arrive at once and are each assigned to a different doctor. Assuming the doctors are assigned randomly to the patients, what is the probability (correct to four decimal places) that all of the assigned physicians are male?
Answer #1

t arsignd phypinrs prob bility that all mab O.0079 aru Acu
Answer #2

We are given here that a clinic employs nine physicians. Five of the physicians are female here.

The probability that none of the assigned 4 physicins are male is computed here as:

= Number of ways to select 4 physicians from 5 female physicians / Total ways to select 4 physicians from 9 physicians

\(=\frac{\left(\begin{array}{l}5 \\ 4\end{array}\right)}{\left(\begin{array}{l}9 \\ 4\end{array}\right)}=0.0397\)

Therefore 0.0397 is the required probability here.

answered by: happy